Arun bought a car for ₹ 3,50,000. The next year, the price went upto ₹ 3,70,000. What was the Percentage of price increase?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question, it is given that
Arun bought a car for =₹350000
The price of the car in the next year, went up to =₹370000
Then increase in price of car =₹370000 -₹350000
=₹20000
The percentage of price increase = (₹20000/₹350000) × 100
= (2/35) × 100
= 200/35
= 40/7
5 whole 5/7

Thanks